# Charactarisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v1

@@ 3. Characterisation of distal actions on the space of one-parameter subgroups
Step 6: Let $\psi,S:H\to H$ be such that $S|_{K}={\rm Id}$, $S|_{C(T)}=\alpha\,{\rm Id}$, $\psi|_{K}={\rm Id}$ and $\psi|_{C(T)}=\phi$, where $\alpha$ and $\phi$ are as in Step 2. Now $T|_{H}=S\psi=\psi S$, and $\psi$ is contained in a compact subgroup of ${\rm Aut}(H)$. Therefore, $S$ acts distally on ${\rm Sub}^{p}_{H}$, and $C(S)=C(T)$. Using that $C(T)=C(S)\neq\{e\}$, we show that $S$ does not act distally on ${\rm Sub}^{p}_{H}$, which would lead to a contradiction.
Let $\{z_{t}=x_{t}c_{t}\}$ be a closed one-parameter subgroup, where $\{x_{t}\}\subset K$ is isomorphic to ${\mathbb{S}}^{1}$ and $\{c_{t}\}\subset C(S)=C(T)$ is a non-trivial (closed) one-parameter subgroup such that $S(c_{t})=c_{\alpha t}$, $t\in{\mathbb{R}}$ (In additive notation on ${\mathbb{R}}^{d}$, we choose $c_{t}=tc$, $t\in{\mathbb{R}}$, for some nonzero $c\in{\mathbb{R}}^{d}$). Let $\{n_{k}\}$ be an unbounded sequence in ${\mathbb{N}}$ such that $S^{n_{k}}(\{z_{t}\})\to C$ for some $C$ in ${\rm Sub}^{p}_{H}$. Note that $\{z_{t}\}\subset\{x_{t}\}\times\{c_{t}\}$ and the latter is $S$-invariant and isomorphic to ${\mathbb{S}}^{1}\times{\mathbb{R}}$. So $C\subset\{x_{t}\}\times\{c_{t}\}$. For a fixed $t\in{\mathbb{R}}$, as $c_{t}\in C(S)=C(T)$, we get that
@@
along a subsequence of $\{n_{k}\}$, for some $s\in{\mathbb{R}}$. Therefore, $x_{s}c_{t}$, and hence, $c_{t}$ belongs to $C$. Thus $\{c_{t}\}\subset C$ and $C=\{x_{t}\}\times\{c_{t}\}$, and hence $C$ is $S$-invariant. This implies that $S^{n_{k}}(S(\{z_{t}\}))\to C$, and hence that $S(\{z_{t}\})=\{z_{t}\}$. As $\{z_{t}\}$ is isomorphic to ${\mathbb{R}}$, there exists $\beta\in{\mathbb{R}}\setminus\{0\}$ such that for all $t\in{\mathbb{R}}$,
@@
As $\{x_{t}\}\times\{c_{t}\}$ is a direct product and $\{c_{t}\}$ is isomorphic to ${\mathbb{R}}$, it follows that $\beta=\alpha$ and $x_{t}=x_{\alpha t}$, i.e. $x_{(1-\alpha)t}=e$ for all $t\in{\mathbb{R}}$. As $\alpha\neq 1$, $x_{t}=e$ for each $t$, a contradiction. Hence, $C(S)=C(T)$ is trivial.
@@
The following corollary which extends Theorem 1.2 is an easy consequence of the theorem in view of Theorem 1.1 of [2]. We will omit the proof.
@@
Since $T$ acts trivially on $G/L$ and $L$, we have that $T$ is distal. In fact, all the eigenvalues of $T$ are equal to 1 and $T$ is unipotent. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Let $U$ be a neighbourhood of the identity $e$ such that every element of $U$ is contained in a one-parameter subgroup. Since $G$ is connected, $U$ generates $G$ and it is enough to show that $T(x)=x$ for all $x\in U$.
Step 1: If possible, suppose $T(x)\neq x$ for some $x\in U$. There exists a one-parameter subgroup $\{x_{t}\}$ in $G$ such that $x_{1}=x$. Let $H=\overline{\{x_{t}\}L}$. Then $H$ is connected solvable and $H/L$ is abelian. Let $\pi:G\to G/L$ be the natural projection. Then $\pi(H)=\overline{\pi(\{x_{t}\})}$. If $\pi(\{x_{t}\})$ is unbounded, it is closed and it is isomorphic to ${\mathbb{R}}$ and in this case $\{x_{t}\}$ is also unbounded and closed. Moreover $H=\{x_{t}\}L$.
@@
Step 2: Let $A:=\{x_{t}\}$ and let $\{n_{k}\}\subset{\mathbb{N}}$ be an unbounded sequence such that $T^{n_{k}}(A)\to B$ for some $B\in{\rm Sub}^{p}_{H}$. By Lemma 2.2, $B$ is not discrete as $A$ is connected and non-trivial. Let $b\in B$ be such that $b\neq e$. There exists $t_{k}\in{\mathbb{R}}$, $k\in{\mathbb{N}}$, such that
@@
Step 3: We first assume that $L$ is simply connected. We show that $y_{t_{0}}=e$ in this case. Let ${\mathfrak{l}}$ be the Lie algebra of $L$. Since $L$ is simply connected and nilpotent, $\exp:{\mathfrak{l}}\to L$ is a homeomorphism with $\log$ as its inverse. Let $y_{t_{k}}=\exp Y_{k}$ for some $Y_{k}\in{\mathfrak{l}}$. Then $y_{t_{k}}^{n_{k}}=\exp(n_{k}Y_{k})$ and as $y_{t_{k}}^{n_{k}}\to y^{\prime}$, we have that $n_{k}Y_{k}\to\log y^{\prime}$ in ${\mathfrak{l}}$. As ${\mathfrak{l}}$ isomorphic to ${\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$, it follows that $Y_{k}\to 0$ in ${\mathfrak{l}}$, and hence $y_{t_{k}}\to e$, i.e. $y_{t_{0}}=e$. Now
@@
Step 4: Suppose that $L$ is not simply connected. Let $M$ be the maximal compact (normal) subgroup of $L$. Then $M$ is central in $G$ and $L/M$ is simply connected. Let $\varrho:G\to G/M$ be the natural projection and let $\overline{T}$ be the automorphism on $G/M$ induced $T$. Then $\overline{T}$ acts trivially on both $\varrho(L)$ and $\varrho(G)/\varrho(L)$. By Proposition 3.4, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$. From Step 3, we get that $\overline{T}$ acts trivially on $G/M$, i.e. $T(g)\in gM$, $g\in G$. In particular, for $\{x_{t}\}$ as above, $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in M$, for all $t\in{\mathbb{R}}$, and $y_{1}\neq e$. Since $M$ is central in $G$, $y_{t}x_{s}=x_{s}y_{t}$, $t,s\in{\mathbb{R}}$, and $\{y_{t}\}$ is a one-parameter subgroup of the compact central subgroup $M$. Note that $\{y_{t}\}$ is closed if $\{x_{t}\}$ is compact.
@@
as $n_{k}\to\infty$. Therefore, $y_{t}\in C$ for each $t$ and $\overline{\{y_{t}\}}\subset C$ as $C$ is closed. Also, $T^{n_{k}}(x_{t})=x_{t}y_{t}^{n_{k}}\to x_{t}y$ along a subsequence of $\{n_{k}\}$ for some $y\in\overline{\{y_{t}\}}\subset C$. Hence, $x_{t}\in C$, $t\in{\mathbb{R}}$. Thus $\{x_{t}\}\times\overline{\{y_{t}\}}\subset C$. Now $T(C)=C$, and we get that $T^{n_{k}}(T(\{x_{t}\}))\to C$. as $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(\{x_{t}\})=(\{x_{t}\})$. Since $T$ is unipotent, we have that $T(x_{t})=x_{t}$ for each $t$. But $T(x_{1})=x_{1}y_{1}\neq x_{1}$. This leads to a contradiction. Therefore, $T(x)=x$ for each $x\in U$, and hence, for each $x\in G$. Thus $T={\rm Id}$. ∎
@@
Step 3: Suppose $G$ is not solvable. Let $R$ be the the radical of $G$. Then as $T$ acts distally on ${\rm Sub}^{p}_{R}$, we have from from Step 2 that $T|_{R}={\rm Id}$.
We first assume that $G$ belongs to class ${\mathcal{C}^{\prime}}$, i.e.  $G$ has no compact subgroup of positive dimension. (For this case, we will follow the proof of Theorem 4.1 in [40] to some extent and use Step 1 instead of Proposition 4.2 of [40]). Now ${\rm Aut}(G)$ is almost algebraic and it has finitely many connected components (see [12]). Thus some power of $T$ belongs to a one-parameter subgroup of ${\rm Aut}(G)^{0}$. Since $T$ is unipotent, we get that $T$ itself belongs to a one-parameter subgroup $\{T_{t}\}$ consisting of unipotent elements as $T=T_{1}$. Also, since $\{T^{n}\}_{n\in{\mathbb{Z}}}$ is co-compact in $\{T_{t}\}$, it is easy to see that each $T_{t}$ also acts distally on ${\rm Sub}^{p}_{G}$.
Let $G^{\prime}=G/R$ and let $T^{\prime}_{t}$ be the automorphism of $G^{\prime}$ induced by $T_{t}$, $t\in{\mathbb{R}}$. Here, $T^{\prime}_{t}$ is also unipotent. Note that $G^{\prime}$ is semisimple and ${\rm Aut}(G^{\prime})^{0}$ is the group of inner automorphisms of $G^{\prime}$. There exists a one-parameter subgroup $\{u_{t}\}$ consisting of Ad-unipotent elements in $G^{\prime}$, such that $T^{\prime}_{t}={\rm inn}(u_{t})$, the inner automorphisms by $u_{t}$, $t\in{\mathbb{R}}$. As shown in [40] we can choose a Iwasawa decomposition $G^{\prime}=KAU$ such that $\{u_{t}\}\subset U$, and $U$ (resp. $AU$) is a closed simply connected nilpotent (resp. solvable) group. Let $\pi:G\to G^{\prime}=G/R$ denote the natural projection. Then $B:=\pi^{-1}(AU)$ is a closed connected solvable group, $T_{t}(B)=B$ and $T_{t}$ acts distally on ${\rm Sub}^{p}_{B}$. This implies from Step 2, that $T_{t}|_{B}={\rm Id}$. In particular, ${\rm Inn}(u_{t})$ acts trivially on $AU$, i.e. $\{u_{t}\}$ is contained in the center of $AU$. Let $\mathcal{A}$ (resp. $\mathcal{U}$) be the Lie algebra of $A$ (resp. $U$). Then $u_{t}=\exp tX$, $t\in{\mathbb{R}}$, for some $X\in\mathcal{U}$, and hence $X$ centralises $\mathcal{A}$. Using the restricted root system and Lemma 6.50 of [28], it follows that the centraliser of $\mathcal{A}$ in $\mathcal{A}\,\mathcal{U}$ is contained in $\mathcal{A}$. Therefore, $X\in\mathcal{A}\cap\mathcal{U}=\{0\}$. This implies that for each $t$, $u_{t}=e$, and hence that $T^{\prime}_{t}={\rm Id}$.
@@
We have shown that under the conditions stated in the proposition, $T={\rm Id}$ if $G$ is either solvable or if $G$ belongs to class ${\mathcal{C}^{\prime}}$.
@@
Let $G$ be a connected Lie group in class ${\mathcal{C}^{\prime}}$. Then ${\rm Aut}(G)$ is almost algebraic as a subgroup of ${\rm GL}({\mathfrak{g}})$ (via the correspondence $T\leftrightarrow{\rm d\,}T$) (cf. [12]). Suppose $T\in{\rm Aut}(G)$ acts distally on ${\rm Sub}^{p}_{G}$. Suppose $G$ is not nilpotent. In particular, it is not a vector group. By Theorem 1.2, $T$ acts distally on $G$. Then for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected abelian almost algebraic subgroup of ${\rm Aut}(G)$ and it has the form ${\mathcal{K}}\times\,{\mathcal{U}}$, where ${\mathcal{K}},\,{\mathcal{U}}\subset{\rm Aut}(G)$, ${\mathcal{K}}$ is an abelian compact connected group and ${\mathcal{U}}$ is an abelian unipotent connected group; i.e. it consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now $T^{n}=S\circ\psi=\psi\circ S$, where $S\in{\mathcal{U}}$ is unipotent and $\psi\in{\mathcal{K}}$. Hence, $S$ acts distally on ${\rm Sub}^{p}_{G}$ (cf. [39], Lemma 2.2). By Theorem 1.3, $S={\rm Id}$. Thus $T^{n}=\psi\in{\mathcal{K}}$ and hence $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse is obvious (cf. [30], Lemma 4.2). ∎
Now we want to prove Theorem 1.4. We need an elementary result about certain automorphisms keeping a maximal compact abelian subgroups (maximal tori) invariant. Maximal compact subgroups in a connected Lie group $G$ are connected and conjugate to each other (see [24]). Any maximal compact subgroup of a connected solvable Lie group is abelian. Any maximal compact abelian subgroup of $G$ is connected (in fact, it is a Cartan subgroup of the maximal compact group in which it is contained). A maximal compact (connected) abelian subgroup of a connected Lie group $G$ is called a maximal torus and maximal tori are conjugate to each other. Any compact abelian subgroup is contained in a maximal torus. If $A$ is a maximal torus of $G$, then so is $T(A)$, for any $T\in{\rm Aut}(G)$, and $T(A)$ is a conjugate of $A$.
In [15], any subgroup of ${\rm Aut}(G)$ which fixes elements of a maximal torus are shown to be almost algebraic. Given any automorphism $T$ of $G$, it need not keep any torus invariant. However, this is the case if $T$ is contained in a compact connected abelian subgroup of ${\rm Aut}(G)$ as shown in the following elementary lemma, which also holds if we replace the radical $R$ by any closed connected characteristic subgroup of $G$.
@@
Let $G$ be a connected Lie group. If ${\mathcal{K}}$ is a compact connected abelian subgroup of ${\rm Aut}(G)$, then there exists a maximal torus of $G$ (resp. of the radical of $G$) which is invariant under every automorphism of $G$ which belongs to ${\mathcal{K}}$.
@@
Let $A$ be any maximal torus in a connected Lie group $G$. For any closed connected normal subgroup $L$ of $G$, since the maximal tori of $G$ (resp. $L$) are conjugate to each other in $G$ (resp. $L$), we get that $A$ contains a maximal torus of $L$ and hence, $A\cap L$ is a maximal torus of $L$.
Let ${\mathcal{K}}$ be a compact connected abelian subgroup of ${\rm Aut}(G)$ and let $H={\mathcal{K}}\ltimes G$; where the action of ${\mathcal{K}}$ on $G$ is by automorphisms. Then $H$ is a connected Lie group and $G$ is normal in $H$. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$ and hence normal in $H$. Let $A$ be a maximal torus in $H$ containing ${\mathcal{K}}$. Arguing as above, we get that $A\cap G$ (resp.  $A\cap R$) is a maximal torus in $G$ (resp. $R$). As $A\cap G$ (resp. $A\cap R$) is normal in $A$ and ${\mathcal{K}}\subset A$, we get that $A\cap G$ (resp. $A\cap R$) is invariant under every automorphism which belongs ${\mathcal{K}}$. ∎
@@
Suppose $G=M$. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], $T^{n}$ is unipotent for some $n\in{\mathbb{N}}$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, so does $T^{n}$, and by Theorem 1.3, $T^{n}={\rm Id}$. Hence the assertion holds when $G$ is a compact and abelian.

### v2

@@ 4. Proofs of the main results
Step 6: Let $\psi,S:H\to H$ be such that $S|_{K}={\rm Id}$, $S|_{C(T)}=\alpha\,{\rm Id}$, $\psi|_{K}={\rm Id}$ and $\psi|_{C(T)}=\phi$, where $\alpha$ and $\phi$ are as in Step 2. Now $T|_{H}=S\psi=\psi S$, and $\psi$ is contained in a compact subgroup of ${\rm Aut}(H)$. Therefore, $S$ acts distally on ${\rm Sub}^{p}_{H}$ (cf. [41], Lemma 2.2), and $C(S)=C(T)$.
Let $\{z_{t}=x_{t}c_{t}\}$ be a closed one-parameter subgroup, where $\{x_{t}\}\subset K$ is isomorphic to ${\mathbb{S}}^{1}$ and $\{c_{t}\}\subset C(S)=C(T)$ is a non-trivial (closed) one-parameter subgroup such that $S(c_{t})=c_{\alpha t}$, $t\in{\mathbb{R}}$ (as addition is the group operation on ${\mathbb{R}}^{d}$, we choose $c_{t}=tc$, $t\in{\mathbb{R}}$, for some nonzero $c\in{\mathbb{R}}^{d}$). Let $\{n_{k}\}$ be an unbounded sequence in ${\mathbb{N}}$ such that $S^{n_{k}}(\{z_{t}\})\to C$ for some $C$ in ${\rm Sub}^{p}_{H}$. Note that $\{z_{t}\}\subset\{x_{t}\}\times\{c_{t}\}$ and the latter is $S$-invariant and isomorphic to ${\mathbb{S}}^{1}\times{\mathbb{R}}$. So $C\subset\{x_{t}\}\times\{c_{t}\}$. For a fixed $t\in{\mathbb{R}}$, as $c_{t}\in C(S)=C(T)$, we get that
@@
along a subsequence of $\{n_{k}\}$, for some $s\in{\mathbb{R}}$. Therefore, $x_{s}c_{t}$, and hence, $c_{t}$ belongs to $C$. Thus $\{c_{t}\}\subset C$ and $C=\{x_{t}\}\times\{c_{t}\}$, and hence $C$ is $S$-invariant. This implies that $S^{n_{k}}(S(\{z_{t}\}))\to C$, and hence that $S(\{z_{t}\})=\{z_{t}\}$ since $S$ acts distally on ${\rm Sub}^{p}_{G}$. As $\{z_{t}\}$ is isomorphic to ${\mathbb{R}}$, there exists $\beta\in{\mathbb{R}}\setminus\{0\}$ such that for all $t\in{\mathbb{R}}$,
@@
As $\{x_{t}\}\times\{c_{t}\}$ is a direct product and $\{c_{t}\}$ is isomorphic to ${\mathbb{R}}$, it follows that $\beta=\alpha$ and $x_{t}=x_{\alpha t}$, i.e. $x_{(1-\alpha)t}=e$ for all $t\in{\mathbb{R}}$. As $\alpha\neq 1$, $x_{t}=e$ for each $t$, a contradiction. Hence, $C(T)$ is trivial.
@@
In view of Theorem 1.1 of [2], the following corollary is an easy consequence of Theorem 1.2. We will omit the proof.
@@
Since $T$ acts trivially on $G/L$ and $L$, we have that $T$ acts distally on $G$. In fact, all the eigenvalues of ${\rm d\,}T$ are equal to 1 and $T$ is unipotent. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Let $U$ be a neighbourhood of the identity $e$ such that every element of $U$ is contained in a one-parameter subgroup. Since $G$ is connected, $U$ generates $G$ and it is enough to show that $T(x)=x$ for all $x\in U$.
Step 1: If possible, suppose $T(x)\neq x$ for some $x\in U$. There exists a one-parameter subgroup $\{x_{t}\}$ in $G$ such that $x_{1}=x$. Let $H=\overline{\{x_{t}\}L}$. Then $H$ is connected and solvable, and $H/L$ is abelian. Let $\pi:G\to G/L$ be the natural projection. Then $\pi(H)=\overline{\pi(\{x_{t}\})}$. If $\pi(\{x_{t}\})$ is unbounded, it is closed and it is isomorphic to ${\mathbb{R}}$ and in this case $\{x_{t}\}$ is also unbounded and closed. Moreover $H=\{x_{t}\}L$.
@@
Step 2: Let $A:=\{x_{t}\}$ and let $\{n_{k}\}\subset{\mathbb{N}}$ be an unbounded sequence such that $T^{n_{k}}(A)\to B$ for some $B\in{\rm Sub}^{p}_{H}$. By Lemma 2.2, $B$ is not discrete as $A$ is connected and non-trivial. Let $b\in B$ be such that $b\neq e$. There exist $t_{k}\in{\mathbb{R}}$, $k\in{\mathbb{N}}$, such that
@@
Step 3: We first assume that $L$ is simply connected. We show that $y_{t_{0}}=e$ in this case. Let ${\mathfrak{l}}$ be the Lie algebra of $L$. Since $L$ is simply connected and nilpotent, $\exp:{\mathfrak{l}}\to L$ is a homeomorphism with $\log$ as its inverse. Let $y_{t_{k}}=\exp Y_{k}$ for some $Y_{k}\in{\mathfrak{l}}$. Then $y_{t_{k}}^{n_{k}}=\exp(n_{k}Y_{k})$ and as $y_{t_{k}}^{n_{k}}\to y^{\prime}$, we have that $n_{k}Y_{k}\to\log y^{\prime}$ in ${\mathfrak{l}}$. As ${\mathfrak{l}}$ is isomorphic to ${\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$, it follows that $Y_{k}\to 0$ in ${\mathfrak{l}}$, and hence $y_{t_{k}}\to e$, i.e. $y_{t_{0}}=e$. Now
@@
Step 4: Suppose that $L$ is not simply connected. Let $M$ be the maximal compact (normal) subgroup of $L$. Then $M$ is central in $G$ and $L/M$ is simply connected. Let $\varrho:G\to G/M$ be the natural projection and let $\overline{T}$ be the automorphism on $G/M$ induced by $T$. Then $\overline{T}$ acts trivially on both $\varrho(L)$ and $\varrho(G)/\varrho(L)$. By Proposition 3.5, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$. From Step 3, we get that $\overline{T}$ acts trivially on $G/M$, i.e. $T(g)\in gM$, $g\in G$. In particular, for $\{x_{t}\}$ as above, $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in M$, for all $t\in{\mathbb{R}}$, and $y_{1}\neq e$. Since $M$ is central in $G$, $y_{t}x_{s}=x_{s}y_{t}$, $t,s\in{\mathbb{R}}$, and $\{y_{t}\}$ is a one-parameter subgroup of the compact central subgroup $M$. Note that $\{y_{t}\}$ is closed if $\{x_{t}\}$ is compact.
@@
as $n_{k}\to\infty$. Therefore, $y_{t}\in C$ for each $t$ and $\overline{\{y_{t}\}}\subset C$ as $C$ is closed. Also, $T^{n_{k}}(x_{t})=x_{t}y_{t}^{n_{k}}\to x_{t}y$ along a subsequence of $\{n_{k}\}$ for some $y\in\overline{\{y_{t}\}}\subset C$. Hence, $x_{t}\in C$, $t\in{\mathbb{R}}$. Thus $\{x_{t}\}\times\overline{\{y_{t}\}}\subset C$. Now $T(C)=C$, and we get that $T^{n_{k}}(T(\{x_{t}\}))\to C$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(\{x_{t}\})=\{x_{t}\}$. Since $T$ is unipotent, we have that $T(x_{t})=x_{t}$ for each $t$. But $T(x_{1})=x_{1}y_{1}\neq x_{1}$. This leads to a contradiction. Therefore, $T(x)=x$ for each $x\in U$, and hence, for each $x\in G$. Thus $T={\rm Id}$. ∎
@@
Step 3: Suppose $G$ is not solvable. Let $R$ be the radical of $G$. Then as $T$ acts distally on ${\rm Sub}^{p}_{R}$, we have from Step 2 that $T|_{R}={\rm Id}$.
We first assume that $G$ belongs to class ${\mathcal{C}^{\prime}}$, i.e.  $G$ has no compact subgroup of positive dimension. (For this case, we will follow the proof of Theorem 4.1 in [42] to some extent and then use Step 1 instead of Proposition 4.2 of [42].) Now ${\rm Aut}(G)$ is almost algebraic and it has finitely many connected components (see [12]). Thus some power of $T$ belongs to a one-parameter subgroup of ${\rm Aut}(G)^{0}$. Since $T$ is unipotent, we get that $T$ itself belongs to a one-parameter subgroup $\{T_{t}\}$ consisting of unipotent elements with $T=T_{1}$. Also, since $\{T^{n}\}_{n\in{\mathbb{Z}}}$ is co-compact in $\{T_{t}\}$, it is easy to see that each $T_{t}$ also acts distally on ${\rm Sub}^{p}_{G}$.
Let $G^{\prime}:=G/R$ and let $T^{\prime}_{t}$ be the automorphism of $G^{\prime}$ induced by $T_{t}$, $t\in{\mathbb{R}}$. Here, $T^{\prime}_{t}$ is also unipotent. Note that $G^{\prime}$ is semisimple and ${\rm Aut}(G^{\prime})^{0}$ is the group of inner automorphisms of $G^{\prime}$. There exists a one-parameter subgroup $\{u_{t}\}$ consisting of Ad-unipotent elements in $G^{\prime}$, such that $T^{\prime}_{t}={\rm inn}(u_{t})$, the inner automorphisms by $u_{t}$, $t\in{\mathbb{R}}$. As shown in [42] we can choose a Iwasawa decomposition $G^{\prime}=KAU$ such that $\{u_{t}\}\subset U$, and $U$ (resp. $AU$) is a closed simply connected nilpotent (resp. solvable) group. Let $\pi:G\to G^{\prime}$ denote the natural projection. Then $B:=\pi^{-1}(AU)$ is a closed connected solvable group, $T_{t}(B)=B$ and $T_{t}$ acts distally on ${\rm Sub}^{p}_{B}$. This implies from Step 2, that $T_{t}|_{B}={\rm Id}$. In particular, ${\rm Inn}(u_{t})$ acts trivially on $AU$, i.e. $\{u_{t}\}$ is contained in the center of $AU$. Let $\mathcal{A}$ (resp. $\mathcal{U}$) be the Lie algebra of $A$ (resp. $U$). Then $u_{t}=\exp tX$, $t\in{\mathbb{R}}$, for some $X\in\mathcal{U}$, and hence $X$ centralises $\mathcal{A}$. Using the restricted root system and Lemma 6.50 of [29], it follows that the centraliser of $\mathcal{A}$ in $\mathcal{A}\,\mathcal{U}$ is contained in $\mathcal{A}$. Therefore, $X\in\mathcal{A}\cap\mathcal{U}=\{0\}$. This implies that for each $t$, $u_{t}=e$, and hence that $T^{\prime}_{t}={\rm Id}$.
@@
We have shown that under the conditions stated in the hypothesis, $T={\rm Id}$ if $G$ is either solvable or if $G$ belongs to class ${\mathcal{C}^{\prime}}$.
@@
Let $G$ be a connected Lie group in class ${\mathcal{C}^{\prime}}$. Then ${\rm Aut}(G)$ is almost algebraic as a subgroup of ${\rm GL}({\mathfrak{g}})$ (via the correspondence $T\leftrightarrow{\rm d\,}T$) (cf. [12]). Suppose $T\in{\rm Aut}(G)$ acts distally on ${\rm Sub}^{p}_{G}$. Suppose $G$ is not nilpotent. In particular, it is not a vector group. By Theorem 1.2, $T$ acts distally on $G$. Then for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected abelian almost algebraic subgroup of ${\rm Aut}(G)$ and it has the form ${\mathcal{K}}\times\,{\mathcal{U}}$, where ${\mathcal{K}},\,{\mathcal{U}}\subset{\rm Aut}(G)$, ${\mathcal{K}}$ is an abelian compact connected group and ${\mathcal{U}}$ is an abelian unipotent connected group; i.e. it consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now $T^{n}=S\circ\psi=\psi\circ S$, where $S\in{\mathcal{U}}$ is unipotent and $\psi\in{\mathcal{K}}$. Hence, $S$ acts distally on ${\rm Sub}^{p}_{G}$ (cf. [41], Lemma 2.2). By Theorem 1.3, $S={\rm Id}$. Thus $T^{n}=\psi\in{\mathcal{K}}$, and hence $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse is obvious (cf. [31], Lemma 4.2). ∎
Now we want to prove Theorem 1.4. We need an elementary result about certain automorphisms keeping a maximal compact connected abelian subgroup (maximal tori) invariant. In a connected Lie group $G$, all maximal tori are conjugate to each other. Any compact abelian subgroup of $G$ is contained in a maximal torus, as any maximal compact abelian subgroup of $G$ is connected. If $A$ is a maximal torus in $G$, then so is $T(A)$, for any $T\in{\rm Aut}(G)$, and $T(A)$ is conjugate to $A$. Any maximal compact subgroup of a connected solvable Lie group is abelian (cf. [25]).
In [15], any subgroup of ${\rm Aut}(G)$ which fixes elements of a maximal torus is shown to be almost algebraic. Given any automorphism $T$ of $G$, it need not keep any torus invariant. However, this is the case if $T$ is contained in a compact connected abelian subgroup of ${\rm Aut}(G)$, as shown in the following elementary lemma. The lemma also holds if we replace the radical $R$ by any closed connected characteristic subgroup of $G$.
@@
Let $G$ be a connected Lie group. If ${\mathcal{K}}$ is a compact connected abelian subgroup of ${\rm Aut}(G)$, then there exists a maximal torus of $G$ (resp. of the radical of $G$) on which every automorphism in ${\mathcal{K}}$ acts trivially.
@@
Let $A$ be any maximal torus in a connected Lie group $G$. It is easy to see that for any closed connected normal subgroup $L$ of $G$, $A\cap L$ is a maximal torus of $L$.
Let ${\mathcal{K}}$ be a compact connected abelian subgroup of ${\rm Aut}(G)$ and let $H={\mathcal{K}}\ltimes G$; where the action of ${\mathcal{K}}$ on $G$ is by automorphisms. Then $H$ is a connected Lie group and $G$ is normal in $H$. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$ and hence normal in $H$. Let $A$ be a maximal torus in $H$ containing ${\mathcal{K}}$. Arguing as above, we get that $A\cap G$ (resp.  $A\cap R$) is a maximal torus in $G$ (resp. $R$). As ${\mathcal{K}}\subset A$ and $A$ s abelian, we get that every automorphism in ${\mathcal{K}}$ acts trivially on $A\cap G$ (resp. $A\cap R$). ∎
@@
Suppose $G=M$. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], $T^{n}$ is unipotent for some $n\in{\mathbb{N}}$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, so does $T^{n}$, and by Theorem 1.3, $T^{n}={\rm Id}$. Hence the assertion holds when $G$ is compact and abelian.